OpenStudy (anonymous):

@blurbendy -5+7b=8b-6

OpenStudy (anonymous):

nvm i know this one

OpenStudy (blurbendy):

-5 + 7b = 8b -6 7b = 8b -1 -b = -1 b = 1 I thought we just did this one, lol

OpenStudy (anonymous):

lol I realized that